# Xuezhang Chen  and Shihong Zhang

### v1

@@
Abstract: We introduce the notion of a biharmonic Poisson kernel associated with certain pair of conformal boundary operators and present its explicit formula. With this powerful tool, we next establish the classification theorems of nonnegative solutions to a biharmonic equation on the upper half-space and unit ball with proper pairs of conformal boundary operators.
$\mathbf{2020}$ MSC: 35B53,53C18 (35G30,35J08)
Keywords: Conformal boundary operators, biharmonic Poisson kernel, biharmonic Green function.
@@ 1 Introduction
Caffarelli-Gidas-Spruck [5] established the celebrated Liouville-type theorem: For $n\geq 2$, a $C^{2}$ solution of
$-\Delta u=(n+1)(n-1)u^{\frac{n+3}{n-1}},~{}~{}u>0\qquad\mathrm{~{}~{}in~{}~{}}\mathbb{R}^{n+1}$
@@
$u(X)=\left(\frac{\varepsilon}{\varepsilon^{2}+|X-X_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},X_{0}\in\mathbb{R}^{n+1}.$
Under an additional hypothesis $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$, the same result was also proved earlier by Obata [28] and Gidas-Ni-Nirenberg [16]. Let $\mathbb{R}_{+}^{n+1}=\{X=(x,t)|~{}x\in\mathbb{R}^{n},t\in\mathbb{R}_{+}\}$ denote the upper half-space with boundary $\partial\mathbb{R}_{+}^{n+1}=\mathbb{R}^{n}$ and $\mathbb{B}^{n+1}$ denote the unit ball in $\mathbb{R}^{n+1}$. Escobar [13] used the Obata-type argument to characterize minimizers of a sharp Sobolev trace inequality on $\mathbb{B}^{n+1}$, via a conformal map between $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$, which is equivalent to seeking smooth solutions to
$\begin{cases}\displaystyle\Delta u=0,\quad u>0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle-\partial_{t}u(x,0)=(n-1)u^{\frac{n+1}{n-1}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.1)
under an extra hypothesis that $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$. Y. Y. Li-Zhu [26] (see also, Y. Y. Li-Zhang [25]) employed the moving sphere method to prove the following Liouville-type theorem without the above hypothesis: A $C^{2}$ solution of (1.1) is of the form
$u(x,t)=\left(\frac{\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
Such Liouville-type theorems of general critical growth second order semilinear elliptic equations on $\mathbb{R}_{+}^{n+1}$ are referred to Escobar [14], Y. Y. Li-Zhang [25] etc.
Sun-Xiong [30] classified nonnegative smooth solutions of the following conformally invariant polyharmonic boundary value problem: For $m\geq 2,n+1>2m$ and $1<p\leq\frac{n+2m-1}{n+1-2m}$,
@@
where $20\leq k\leq m-$ 2. When $m$ is even, assume in addition that $u(X)=o(|X|^{2m-1})$ as $|X|\to\infty$, then for the critical exponent $p=\frac{n+2m-1}{n-(2m-1)}$,
$u(x,t)=$ $c(n,m)\int_{\mathbb{R}^{n}}\frac{t^{2m-1}}{(|x-y|^{2}+t^{2})^{\frac{n+2m-1}{2}}}\left(\frac{\lambda}{\lambda^{2}+|y|^{2}}\right)^{\frac{n-2m+1}{2}}\mathrm{d}y+\sum_{k=1}^{m-1}t^{2k}P_{2k}(x),$
where $x_{0}\in\mathbb{R}^{n},\lambda\in\mathbb{R}_{+}$, $c(n,m)$ is a normalized constant and $P_{2k}(x)$ denotes a polynomial of degree $2k$ in the variable $x$; for $1<p<\frac{n+2m-1}{n+1-2m}$,
$u(x,t)=\sum_{k=1}^{m-1}t^{2k}P_{2k}(x).$
A key step of their approach is to improve the integrability of a Kevin transformation $u^{\ast}$ for $(-\Delta)^{m}$ via an iteration method, and eventually to show that the singularity of $u^{\ast}$ can be removed at $t=0$; in the next step, the integral equation of $u^{\ast}$ on $\partial\mathbb{R}_{+}^{n+1}$ is solved by the classification theorems due to Chen-Li-Ou [9] and Y. Y. Li [24], then the standard method can give the classification of nonnegative solutions to (1.2).
Conformal boundary operators $\mathscr{B}_{k}^{3}$ for $k\in\{0,1,2,3\}$ associated with the Paneitz operator have attracted a lot of attention, see [18, 4, 6, 11] etc. In the study of zeta functional determinants on a four-manifold with boundary, Chang-Qing [8] discovered a third order conformally covariant boundary operator $\mathscr{B}_{3}^{3}$. The discovery of conformal boundary operators $\mathscr{B}_{k}^{3}$ to an $(n+1)$-manifold $(X^{n+1},g)$ with boundary $\partial X^{n+1}=M^{n}$ is due to Branson-Gover [3], Juhl [23], Gover-Peterson [19], Grant [20], Stafford [29] etc. The Paneitz operator $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ enjoy the property (cf. [7, Theorem 1.1]) that if $\tilde{g}=e^{2\tau}g$, then for all $\Psi\in C^{\infty}(\overline{X^{n+1}})$,
@@
Readers are referred to Case [7] or authors [11] for explicit formulae of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ on manifolds with boundary. We are concerned with two conformally equivalent model spaces: $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$. On the upper half-space (+$(\mathbb{R}^{n+1}_{+},\partial\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})$, the explicit formulae of conformal boundary operators are $\mathscr{B}^{3}_{0}(u)=u$ and
$\mathscr{B}^{3}_{1}(u)=$ $-\partial_{t}u;$
$\mathscr{B}^{3}_{2}(u)=$ $\partial_{t}^{2}u-\overline{\Delta}u;$
$\mathscr{B}^{3}_{3}(u)=$ $\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u$
for all $u\in C^{\infty}(\overline{\mathbb{R}_{+}^{n+1}})$. On the unit ball $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, these conformal boundary operators are $\mathscr{B}^{3}_{0}(U)=U$ and
@@
for all $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, where $r=|\xi|$ for $\xi\in\overline{\mathbb{B}^{n+1}}$; the corresponding boundary curvatures $(T_{i}^{3})_{g_{\mathbb{S}^{n}}}$ are constant and set $\mathbb{T}^{3}_{i}=\frac{n-3}{2}(T_{i}^{3})_{g_{\mathbb{S}^{n}}}$ for $1\leq i\leq 3$, explicitly,
@@
Let $\Delta$ and $\overline{\Delta}$ denote the Euclidean Laplacians on $\mathbb{R}^{n+1}$ and $\mathbb{R}^{n}$, respectively. Let $p_{k}^{\ast}=\frac{n+2k-3}{n-3}$ for $k\in\{0,1,2,3\}$, as $p_{k}^{\ast}+1$ is the critical Sobolev exponent associated to $\mathscr{B}^{3}_{k}$.
This motivates us to classify nonnegative smooth solutions to biharmonic boundary value problems on $\mathbb{R}_{+}^{n+1}$ associated with pairs of conformal boundary operators $(\mathscr{B}^{3}_{i},\mathscr{B}^{3}_{j})$ for $i<j\in\{1,2,3\}$: For $p_{i},p_{j}>0$, provided $u(x,0)$ has some suitable decay at infinity,
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{i}u^{p_{i}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(u)=\mathbb{T}^{3}_{j}u^{p_{j}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.5)
The nonlinear interaction among $\mathscr{B}^{3}_{i}$ and $\mathscr{B}^{3}_{j}$ brings us new obstacles, compared with vanishing lower order part of derivatives in [30]. So, it has great difficulty in applying the approach in [30] to (1.5), as well as some peculiar phenomena occur.
An extremely powerful tool in our approach is so-called biharmonic Poisson kernel consisting of two kernel functions corresponding to a proper pair of conformal boundary operators $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$. In the following, we separate our discussion into two cases: $n\neq 3$ and $n=3$. In the latter case, it involves the critical dimension for both the Paneitz operator and $\mathscr{B}_{3}^{3}$.
Definition 1.1 (Poisson kernel for a biharmonic boundary value problem).
For $i,j\in\{0,1,2,3\}$ and $i<j$, a pair of kernel functions $P_{i}^{3}\oplus P_{j}^{3}(X)$ for $X\in\mathbb{R}_{+}^{n+1}$ is called a biharmonic Poisson kernel associated to a pair of conformal boundary operators $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$, if the following Poisson integral
@@
$\begin{cases}\displaystyle\Delta^{2}v=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(v)=f_{i}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(v)=f_{j}\qquad\,\,\,\,\,\,\,&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.6)
Obviously, the notion of biharmonic Poisson kernel could be naturally extended to $\mathbb{B}^{n+1}$ with boundary condition $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$.
@@
For $n\neq 3$ and $X=(x,t)\in\mathbb{R}_{+}^{n+1}$, we let
@@
$P_{3}^{3}(X)=$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{(t^{2}+|x|^{2})^{\frac{n-3}{2}}}.$
Suppose (i) $f_{0}\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ ; (ii) $k\in\{1,2,3\}$, $f_{k}\in C^{2}(\mathbb{R}^{n})$ and $f_{k}=O(|x|^{-k-\delta_{k}})$ as $|x|\to\infty$ for some $\delta_{k}\in\mathbb{R}_{+}$. Then for $i,j\in\{0,1,2,3\}$ and $i+j\neq 3$, the explicit formula of Poisson kernel corresponding to biharmonic boundary value problem (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$.
@@
It is straightforward to verify that other than two pairs $(\mathscr{B}_{0}^{3},\mathscr{B}_{3}^{3})$ and $(\mathscr{B}_{1}^{3},\mathscr{B}_{2}^{3})$, each $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$ satisfies ‘the complementing condition’ introduced by Agmon-Douglis-Nirenberg [1, Section 10], see also [15, Section 2.3].
For a smooth compact Riemannian manifold $(X^{n+1},g)$ with boundary $M^{n}$ and $n\geq 4$, under the assumption that a certain conformal invariant $\lambda_{1}(P_{4}^{g})>0$ (see [7]), the existence and uniqueness of smooth solutions to the following boundary value problem was studied in [7, Proposition 1.3]: For $\bar{f}_{0},\bar{f}_{1}\in C^{\infty}(M^{n})$,
$\begin{cases}\displaystyle P_{4}^{g}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad X^{n+1},\\ \displaystyle\mathscr{B}^{3}_{0}U=\bar{f}_{0}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,M^{n},\\ \displaystyle\mathscr{B}^{3}_{1}U=\bar{f}_{1}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,M^{n}.\end{cases}$
On the model space $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, for each proper pair of conformal boundary operators, such a biharmonic boundary value problem is expected to be better understood, serving as a first step to the prescribing boundary curvature problem on manifolds with boundary.

### v2

@@
Abstract: For conformal boundary operators associated with the Paneitz operator, we introduce a rigorous definition of the biharmonic Poisson kernel consisting of a pair of kernel functions and derive its explicit representation formula. With this powerful tool, we establish classification theorems of nonnegative solutions to the $Q$-flat and constant $T$-curvature equations on $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
$\mathbf{2020}$ MSC: 35J40, 53C18 (35J91, 53C24)
Keywords: $Q$-curvature, conformal boundary operator, $T$-curvature, biharmonic Poisson kernel.
@@ 1 Introduction
In a seminal work [7], Caffarelli-Gidas-Spruck established a Liouville-type theorem of the second-order critical semilinear elliptic equation: For $n\geq 2$, every positive $C^{2}$ solution to
$-\Delta u=(n+1)(n-1)u^{\frac{n+3}{n-1}},\qquad\quad\mathrm{~{}~{}in~{}~{}}\mathbb{R}^{n+1}$
@@
$u(X)=\left(\frac{\varepsilon}{\varepsilon^{2}+|X-X_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\mathrm{~{}~{}for~{}~{}some~{}~{}}\varepsilon\in\mathbb{R}_{+},X_{0}\in\mathbb{R}^{n+1}.$
Under an additional hypothesis $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$, the same result was also proved earlier by Obata [34] and Gidas-Ni-Nirenberg [21]. Escobar [17] used the Obata-type argument to characterize minimizers of a sharp Sobolev trace inequality on the unit ball $\mathbb{B}^{n+1}$, via a conformal map between $\mathbb{R}_{+}^{n+1}:=\{X=(x,t)\in\mathbb{R}^{n}\times\mathbb{R}_{+}\}$ and $\mathbb{B}^{n+1}$, which is equivalent to seeking positive smooth solutions to
$\begin{cases}\displaystyle\Delta u=0,&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle-\partial_{t}u(x,0)=(n-1)u^{\frac{n+1}{n-1}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.1)
under a decay hypothesis that $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$. Y. Y. Li-Zhu [32] (see also, Y. Y. Li-Zhang [31]) employed the moving sphere method to prove the following Liouville-type theorem without the decay hypothesis: Every positive $C^{2}$ solution to (1.1) is of the form
$u(x,t)=\left(\frac{\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\mathrm{~{}~{}for~{}~{}some~{}~{}}\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
Such Liouville-type theorems of critical semilinear second-order elliptic equations on $\mathbb{R}_{+}^{n+1}$ are referred to Escobar [18], Chipot-Shafrir-Fila [11], Y. Y. Li-Zhang [31] and references therein.
Sun-Xiong [36] classified nonnegative smooth solutions of the following conformally invariant polyharmonic boundary value problem (BVP): For $m\geq 2,n+1>2m$ and $1<p\leq\frac{n+2m-1}{n+1-2m}$,
@@
where $20\leq k\leq m-$ 2. When $m$ is even, assume in addition that $u(X)=o(|X|^{2m-1})$ as $|X|\to\infty$. Then for the critical Sobolev exponent $p=\frac{n+2m-1}{n-(2m-1)}$,
$u(x,t)=$ $c(n,m)\int_{\mathbb{R}^{n}}\frac{t^{2m-1}}{(|x-y|^{2}+t^{2})^{\frac{n+2m-1}{2}}}\left(\frac{\lambda}{\lambda^{2}+|y|^{2}}\right)^{\frac{n-2m+1}{2}}\mathrm{d}y+\sum_{k=1}^{m-1}t^{2k}P_{2m-2-2k}(x),$
where $x_{0}\in\mathbb{R}^{n},\lambda\in\mathbb{R}_{+}$, $c(n,m)$ is a normalized constant and $P_{2m-2-2k}(x)$ denotes a polynomial of degree $\leq 2m-2-2k$ satisfying $\liminf_{|x|\to\infty}P_{2m-2-2k}(x)\geq 0$; for $1<p<\frac{n+2m-1}{n+1-2m}$,
$u(x,t)=\sum_{k=1}^{m-1}t^{2k}P_{2m-2-2k}(x),$
where $P_{2m-2-2k}$ is a nonnegative polynomial of degree $\leq 2m-2-2k$. One of crucial steps in [36] is to improve the integrability of a Kevin transformation $u^{\ast}$ for $(-\Delta)^{m}$ through an iteration method, and to show that the singularity of $u^{\ast}$ can be eventually removed at $t=0$. Next the integral equation of $u^{\ast}$ on $\partial\mathbb{R}_{+}^{n+1}$ is solved by the classification theorems due to Y. Y. Li [30] and Chen-Li-Ou [12], then the classification of solutions to the BVP (1.2) follows by the the standard procedure.
Conformal boundary operators $\mathscr{B}_{k}^{3}$ for $k\in\{1,2,3\}$ associated with the Paneitz operator have attracted a lot of attention, see [23, 6, 8, 15] etc. On a smooth four-manifold with boundary, Chang-Qing [10] discovered a third-order conformal boundary operator $\mathscr{B}_{3}^{3}$ and its associated curvature $T_{3}^{3}$. To generalize the Chang-Qing boundary operator to an $(n+1)$-manifold $(X^{n+1},g)$ with boundary $\partial X:=M^{n}$ for all conformal boundary operators $\mathscr{B}_{k}^{3},k\in\{1,2,3\}$, Branson-Gover [5] initiated the Dirichlet-to-Neumann programme for all the higher order GJMS operators, and introduced the first- and second-order conformal boundary operators $\mathscr{B}_{k}^{3}$ associated with the Paneitz operator. Grant [25] and Stafford [35] made suitable adjustments of the approach in [5] to conduct $\mathscr{B}_{3}^{3}$. Readers are referred to Juhl [29], Gover-Peterson [24] for other treatments. The Paneitz operator $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ enjoy the property that if $\tilde{g}=e^{2\tau}g$, then for all $\Psi\in C^{\infty}(\overline{X^{n+1}})$,
@@
For future reference, adopting the notations as in [10] and [24], we call $(T_{k})_{g}:=\frac{2}{n-3}\mathscr{B}_{k}^{3}(1),k\in\{1,2,3\}$ the $T$-curvatures on $M$, as the higher order generalization of the boundary mean curvature $h_{g}=\frac{2}{n-3}\mathscr{B}_{1}^{3}(1)$. Readers are referred to Case [9] or the authors [15] for explicit formulae of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ on manifolds with boundary.
We are concerned with two models: $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$. On (+$(\mathbb{R}^{n+1}_{+},\partial\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})$, the explicit formulae of conformal boundary operators are $\mathscr{B}^{3}_{0}(u)=u$ and
$\mathscr{B}^{3}_{1}(u)=-\partial_{t}u;\qquad\mathscr{B}^{3}_{2}(u)=\partial_{t}^{2}u-\overline{\Delta}u;\qquad\mathscr{B}^{3}_{3}(u)=$ $\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u$
for all $u\in C^{\infty}(\overline{\mathbb{R}_{+}^{n+1}})$. On $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, these conformal boundary operators are $\mathscr{B}^{3}_{0}(U)=U$ and
@@
for all $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, where $r=|\xi|$ for $\xi\in\overline{\mathbb{B}^{n+1}}$; the corresponding $T$-curvatures $(T_{i}^{3})_{\mathbb{S}^{n}}$ are constant and set $\mathbb{T}^{3}_{i}=\frac{n-3}{2}(T_{i}^{3})_{\mathbb{S}^{n}}$ for $i\in\{1,2,3\}$, explicitly,
@@
Let $\Delta$ and $\overline{\Delta}$ denote Euclidean Laplacians on $\mathbb{R}^{n+1}$ and $\mathbb{R}^{n}$, respectively. Let $p_{k}^{\ast}=\frac{n+2k-3}{n-3}$ for $k\in\{1,2,3\}$, as $p_{k}^{\ast}+1$ is the critical Sobolev exponent associated to $\mathscr{B}^{3}_{k}$.
We are interested in the classification of nonnegative solutions to the $Q$-flat and constant $T$-curvature type equations on $\mathbb{R}_{+}^{n+1}$: Let $n\geq 4$ and $p_{i},p_{j}>0$ for $i,j\in\{1,2,3\},i<j$, consider nonnegative smooth solutions to
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{i}u^{p_{i}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(u)=\mathbb{T}^{3}_{j}u^{p_{j}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.5)
under hypothesis that $u(x,0)$ has some suitable decay at infinity. The nonlinear interaction among $\mathscr{B}^{3}_{i}$ and $\mathscr{B}^{3}_{j}$ brings us new obstacles, compared with vanishing of lower order derivative boundary conditions in [36]. So, it has great difficulty in applying the approach in [36] to (1.5), and peculiar phenomena should occur simultaneously.
Although the Poisson kernel of a higher order elliptic BVP might appear elsewhere in the literature, a rigorous definition of the biharmonic Poisson kernel to the fourth-order elliptic BVP (1.5) is still in demand.
Definition 1.1 (Biharmonic Poisson kernel).
For $i,j\in\{0,1,2,3\},i<j$, a pair of kernel functions $P_{i}^{3}\oplus P_{j}^{3}(X)$ for $X\in\mathbb{R}_{+}^{n+1}$ is called a biharmonic Poisson kernel associated to a conformal boundary operator pair $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$, if the following Poisson integral
@@
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}v=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(v)=f_{i}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(v)=f_{j}\qquad\,\,\,\,\,\,\,&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.6)
We are now ready to give the explicit formulae of biharmonic Poisson kernels.
@@
For $n\geq 2$ and $X=(x,t)\in\mathbb{R}_{+}^{n+1}$, we let111 A heuristic observation allows us to find these explicit kernel functions $P_{k}^{3},k\in\{0,1,2,3\}$.
@@
and
$P_{3}^{3}(X)=\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{(t^{2}+|x|^{2})^{\frac{n-3}{2}}},\quad$ $\qquad\mathrm{~{}~{}when~{}~{}}n\neq 3;$
$P_{3}^{3}(X)=-\frac{1}{4|\mathbb{S}^{3}|}\log(t^{2}+|x|^{2})+C,\quad C\in\mathbb{R},$ $\qquad\mathrm{~{}~{}when~{}~{}}n=3.$
Suppose (i) $f_{0}\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ ; (ii) $k\in\{1,2,3\}$, $f_{k}\in C^{2}(\mathbb{R}^{n})$ and $f_{k}=O(|x|^{-k-\delta_{k}})$ as $|x|\to\infty$ for some $\delta_{k}\in\mathbb{R}_{+}$. Then for $i,j\in\{0,1,2,3\},i+j\neq 3$, the explicit formula of Poisson kernel corresponding to problem (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$.
@@
It is direct to verify that other than $(\mathscr{B}_{0}^{3},\mathscr{B}_{3}^{3})$ and $(\mathscr{B}_{1}^{3},\mathscr{B}_{2}^{3})$, the BVP $(\Delta^{2};\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$ for problem (1.6) satisfies the Lopatinski-Shapiro condition (e.g., Branson-Gover [5, Section 6]; or [9, p.311]) or complementing condition (e.g., Agmon-Douglis-Nirenberg [1, Section 10] or [20, Section 2.3]).
Let
$\Gamma(X-Y)=\Gamma(|X-Y|)=\begin{cases}\displaystyle\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|X-Y|^{3-n}\qquad&\mathrm{~{}~{}when~{}~{}}n\neq 3,\\ \displaystyle-\frac{1}{4|\mathbb{S}^{3}|}\log|X-Y|&\mathrm{~{}~{}when~{}~{}}n=3.\end{cases}$
denote the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$.
Remark 1.2.
As a traditional way of the classical Poisson kernel, conformal boundary operators $\mathscr{B}_{k}^{3}$ indeed act as a bridge between the biharmonic Poisson kernel and fundamental solution: Let $n\geq 2$, for every fixed $X=(x,t)\in\mathbb{R}_{+}^{n+1}$ and all $Y=(y,0)\in\partial\mathbb{R}_{+}^{n+1}$, a direct computation yields
$\begin{split}\mathscr{B}^{3}_{0}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{3}(x-y,t);&\qquad\mathscr{B}^{3}_{1}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{2}(x-y,t);\\ \mathscr{B}^{3}_{2}(\Gamma(X-Y))=-\frac{1}{2}P^{3}_{1}(x-y,t);&\qquad\mathscr{B}^{3}_{3}(\Gamma(X-Y))=-\frac{1}{2}P^{3}_{0}(x-y,t).\end{split}$ (1.7)
Consider
$\begin{cases}\displaystyle\quad\Delta^{2}v=f&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(v)=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(v)=0\qquad\,\,\,\,\,\,\,&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.8)
Definition 1.2 (Biharmonic Green function).
Let $n\geq 3$, $i,j\in\{0,1,2,3\},i+j\neq 3$ and $f\in C^{\infty}(\mathbb{R}_{+}^{n+1})$. A function $G^{(i,j)}:(X,Y)\in\overline{\mathbb{R}_{+}^{n+1}}\times\overline{\mathbb{R}_{+}^{n+1}}\mapsto\mathbb{R}\cup\{\infty\}$ for (1.8) is called a Green function, if $G^{(i,j)}(X,Y)=\Gamma(X-Y)+H^{(i,j)}(X,Y)$, where $H^{(i,j)}$ is a solution of
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}_{+}^{n+1}.\end{cases}$ (1.9)
Remark 1.3.
As a direct consequence of Theorem 1.1 and (1.7), problem (1.9) admits a unique classical solution in $\mathbb{R}_{+}^{n+1}$ for $n\geq 3$.
More importantly, we can derive the explicit formulae of $H^{(i,j)}$ with the help of the Boggio formula in Boggio [4] or Gazzola-Grunau-Sweers [20, Lemma 2.27].
For clarity, we distinguish these representation formulae into two cases: noncritical and critical dimensions.